Classify each of the following as a Lewis acid or a Lewis base.

Drag the appropriate items to their respective bins.

LEWIS ACID OR LEWIS BASE'

1) CO2

2)P(CH3)3

3)H2O

4)B(CH3)3

5)FE3+

6)CN-

7)OH-

8)H+
Answer #1
Concepts and reason

Any chemical substance can be classified as an acid or base by using Lewis acid base theory. Further,the acidic and basic nature of the substance can be analyzed through the ability of molecules to accept and donate electrons.

Fundamentals

According to Lewis acid -base theory,

An acid is a substance that accepts a pair of electrons.

A base is a substance that donates a pair of electrons.

The given molecules are

CO2{\rm{C}}{{\rm{O}}_{\rm{2}}} , P(CH3)3{\rm{P}}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{3}}} , H2O{{\rm{H}}_{\rm{2}}}{\rm{O}} , and B(CH3)3{\rm{B}}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{3}}}

CO2{\rm{C}}{{\rm{O}}_{\rm{2}}} is the electron pair acceptor.

P(CH3)3{\rm{P}}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{3}}} is the electron pair donor.

H2O{{\rm{H}}_{\rm{2}}}{\rm{O}} can act as both electron pair acceptor and donor.

B(CH3)3{\rm{B}}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{3}}} is the electron pair acceptor

The given molecules are Fe3+{\rm{F}}{{\rm{e}}^{{\rm{3 + }}}} , CN{\rm{C}}{{\rm{N}}^{\rm{ - }}} , OH{\rm{O}}{{\rm{H}}^{\rm{ - }}} , H+{{\rm{H}}^{\rm{ + }}} .

Fe3+{\rm{F}}{{\rm{e}}^{{\rm{3 + }}}} is the electron pair acceptor.

CN{\rm{C}}{{\rm{N}}^{\rm{ - }}} is the electron pair donor.

OH{\rm{O}}{{\rm{H}}^{\rm{ - }}} is the electron pair donor.

H+{{\rm{H}}^{\rm{ + }}} is the electron pair acceptor.

Ans:

The molecules that are Lewis acids are CO2{\rm{C}}{{\rm{O}}_{\rm{2}}} , H+{{\rm{H}}^{\rm{ + }}} , H2O{{\rm{H}}_{\rm{2}}}{\rm{O}} , Fe3+{\rm{F}}{{\rm{e}}^{{\rm{3 + }}}} and B(CH3)3{\rm{B}}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{3}}} .

The molecules that are Lewis bases are OH{\rm{O}}{{\rm{H}}^{\rm{ - }}} , CN{\rm{C}}{{\rm{N}}^{\rm{ - }}} H2O{{\rm{H}}_{\rm{2}}}{\rm{O}} and P(CH3)3{\rm{P}}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{3}}} .